OpenStudy (anonymous):

Simplify \[\frac{a^3-b^-3}{b-a}*frac{6}{2a^+2ab+b^2}\]

mathslover (mathslover):

please re write the problem

OpenStudy (anonymous):

\[\frac{a^3-b^3}{b-a}*\frac{6}{2a^2+2ab+b^2}\] I got 3

mathslover (mathslover):

k wait

OpenStudy (anonymous):

Ugh, sorry correct \[\frac{a^3-b^3}{b-a}*\frac{6}{2a^2+2ab+2b^2}\]

mathslover (mathslover):

the correct answer is -3

OpenStudy (anonymous):

hold up

OpenStudy (anonymous):

has to be a minus in the right side, since when you multiply out the right side you need things to cancel so there must be OPPOSITE signs in the first factor (a+b) and the second factor (a^2-ab+b^2) to make that cancellation happen. Similarly, in a^3-b^3=(a-b)(a^2+ab+b^2), you have to have OPPOSITE signs in the two factors on the right side or there will be no cancellation. witch =3

mathslover (mathslover):

u have to use integration here

OpenStudy (anonymous):

yes yes

OpenStudy (anonymous):

ok @truck I don't understand what you're saying... and @mathslover ingeration? This is for algebra I

OpenStudy (anonymous):

\[\frac{a^3-b^3}{b-a}*\frac{6}{2a^2+2ab+2b^2}\]\[=\frac{(a-b)(a^2 + ab+b^2)}{b-a}*\frac{6}{2(a^2+ab+b^2)}\]\[=-1*3\]\[=-3\]

mathslover (mathslover):